Examine whether a best critical region exists for testing the null hypothesis H0: θ=θ0 against the alternative hypothesis H1:θ>θ0 for the parameter θ of the distribution:


f(X, θ)=(1+θ)/(x+2)^2 1≤x<∞
1
Expert's answer
2020-03-06T12:08:09-0500

Suppose "x_1,\ldots,x_n" is a random sample from a population. We will find the test with the best critical region, that is, find the uniformly most powerful test, with a sample size "n" and a significance level "\alpha" to test the simple "H_0: \theta=\theta_0" against the composite "H_1: \theta>\theta_0".

Let "\alpha=P(C;\theta_0)".

For each simple "H_1: \theta=\theta_a", say, the ratio of the likelihood functions is:

"\frac{L(\theta_0)}{L(\theta_a)}=\frac{\frac{1+\theta_0}{(x_1+2)^2}\ldots\frac{1+\theta_0}{(x_n+2)^2}}{\frac{1+\theta_a}{(x_1+2)^2}\ldots\frac{1+\theta_a}{(x_n+2)^2}}\leq k, k=const.\\
(\frac{1+\theta_0}{1+\theta_a})^n\leq k.\\
\frac{1+\theta_0}{1+\theta_a}\leq k^{1/n}."

Here we use the Neyman Pearson Lemma.

The left side of the last inequality does not depend on "x_1,\ldots,x_n". So each critical region is "the best". The best critical region does not exist.